OpenStudy (anonymous):

A CAR TRAVELS 120 miles from A th B at 30 mph but returns the same distance at 40 mph. What is the average speed for the round trip? Please show steps @Algebra I

OpenStudy (agreene):

\[\frac{120mi}{30mph}=4hr\] \[\frac{120mi}{40mph}=3hr\] \[4hr+3hr=7hr\] \[\frac{120mi+120mi}{7hr}=\frac{240mi}{7hr}\approx 34.2857\frac{mi}{hr}\] about 34.2857 mph

OpenStudy (anonymous):

The total you have traveled 240 miles. you travelled 120 miles at 40 mph in t1. v=s/t t1=120/40 =3hours you travelled 120 miles at 30 mph in t2. t2=120/30 =4 hours so your average speed was total distance over total time spent. Thus V_avrg = 240/(t1 +t2) = 240/7 = 34,2857143 mph